$SPAD/src/input polygamma.input

Timothy Daly December 3, 2016

Abstract

Contents

2

— * —

)set break resume
)sys rm -f polygamma.output
)spool polygamma.output
)set message test on
)set message auto off
)clear all
--S 1 of 6
sum(1/k^2,k=1..n)
--R
--R         n
--R        --+    1
--R   (1)  >     --
--R        --+    2
--R        k= 1  k
--R                                         Type: Union(Expression(Integer),...)
--E 1
--S 2 of 6
f1(n)==sum(1/k^2, k=1..n)
--R                                                                   Type: Void
--E 2
--S 3 of 6
f2(n)==-polygamma(1,n+1)+1/6*%pi^2
--R                                                                   Type: Void
--E 3
--S 4 of 6
f2(5)
--R   Compiling function f2 with type PositiveInteger -> DoubleFloat
--R
--R   (4)  1.463611111111111
--R                                                            Type: DoubleFloat
--E 4
--S 5 of 6
f1(5)
--R   Compiling function f1 with type PositiveInteger -> Union(Fraction(
--R      Polynomial(Integer)),Expression(Integer))
--R
--R        5269
--R   (5)  ----
--R        3600
--R                                         Type: Union(Expression(Integer),...)
--E 5

--S 6 of 6
%::DFLOAT
--R
--R   (6)  1.4636111111111112
--R                                                            Type: DoubleFloat
--E 6
)spool
)lisp (bye)

———-

References^{[1] nothing}

5